# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 6 Approach to Chemical Equilibrium
$\mathcal{R}_{r}=f_{k}^{\mathrm{NB}}\,k_{d}^{\mathrm{arr}}(T)\,\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)\left[Z(T,T_{v})\left(\frac{\rho_{\mathrm{AB}}}{M_{\mathrm{AB}}}\right)-\frac{1}{K_{r}^{eq}(T)}\,\left(\frac{\rho_{\mathrm{A}}}{M_{\mathrm{A}}}\right)\left(\frac{\rho_{\mathrm{B}}}{M_{\mathrm{B}}}\right)\right].$ (25)
This expression makes it clear that the non-Boltzmann factor re-scales both forward and backward rates simultaneously. When cold freestream gas initially traverses the shock front, only the forward reaction will contribute to Eq. (25) and the model accurately reproduces $\mathcal{R}_{r}$ during the early QSS dissociation phase. Then, as the gas is advected downstream, vibrational and translation-rotational modes undergo relaxation and the mixture gradually approaches thermo-chemical equilibrium. Forward and backward reactions should then approach their thermal equilibrium rates, i.e. $\mathcal{R}_{r}^{f,\,eq}=k_{d}^{\mathrm{arr}}(T)\,(\rho_{\mathrm{AB}}^{eq}/M_{\mathrm{AB}})(\rho_{\mathrm{M}}/M_{\mathrm{M}})$ and $\mathcal{R}_{r}^{b,\,eq}=k_{d}^{\mathrm{arr}}(T)/K_{r}^{eq}\,(\rho_{\mathrm{A}}^{eq}/M_{\mathrm{A}})(\rho_{\mathrm{B}}^{eq}/M_{\mathrm{B}})(\rho_{\mathrm{M}}/M_{\mathrm{M}})$ respectively. However, when a constant non-Boltzmann factor $f_{k}^{\mathrm{NB}}$ is employed in evaluating the nonequilibrium forward and backward rate coefficients, it will never disappear and approaching these thermal rates becomes impossible.
@@
A related issue arises for the vibrational energy-chemistry transfer term. Evaluating Eq. (8) together with Eqs. (10) and (17) for the present reaction yields:
@@
$Q_{v,r,\mathrm{AB}}^{\mathrm{chem}}=-f_{k}^{\mathrm{NB}}\,f_{\varepsilon}^{\mathrm{NB}}\,k_{d}^{\mathrm{arr}}(T)\,\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)\left[Z(T,T_{v})\left(\frac{\rho_{\mathrm{AB}}}{M_{\mathrm{AB}}}\right)-\frac{1}{K_{r}^{eq}(T)}\,\left(\frac{\rho_{\mathrm{A}}}{M_{\mathrm{A}}}\right)\left(\frac{\rho_{\mathrm{B}}}{M_{\mathrm{B}}}\right)\right]\,\langle\varepsilon_{v,\mathrm{AB}}\rangle_{r}^{\mathrm{Knab}}(T,T_{v}),$ (26)
@@
with both non-Boltzmann factors seen to be re-scaling the coupling source term. As was the case for the reaction rates themselves, this does not constitute a problem during the early QSS dissociation phase, where only the forward-reaction portion contributes meaningfully to $Q_{v,r,\mathrm{AB}}^{\mathrm{chem}}$. However, when constant Boltzmann factors are employed they continue to incorrectly re-scale both forward and reverse vibrational energy-chemistry removal/replenishment rates even as the gas approaches thermal and chemical equilibrium.
@@
To address this problem, we use variable non-Boltzmann factors $f_{k}^{\mathrm{VNB}}$ and $f_{\varepsilon}^{\mathrm{VNB}}$, whose values change dynamically with the local state of the gas. Following the approach of Singh and Schwartzentruber [53] we use the reaction’s nonequilibrium concentration ratio:
$z_{r}=\frac{(\rho_{\mathrm{A}}/M_{\mathrm{A}})\,(\rho_{\mathrm{B}}/M_{\mathrm{B}})}{(\rho_{\mathrm{AB}}/M_{\mathrm{AB}})\,K_{r}^{eq}(T)},$ (27)
@@
which will depend on the local mixture composition and the equilibrium constant for reaction $r$ at temperature $T$. Three limiting cases are of particular interest. (1) When the gas is close to “fully recombined”, i.e. $\rho_{\mathrm{A}}\rightarrow 0$ and $\rho_{\mathrm{B}}\rightarrow 0$, so $z_{r}\rightarrow 0$. (2) When the gas takes on its equilibrium composition $\rho_{\mathrm{A}}=\rho_{\mathrm{A}}^{eq}$, $\rho_{\mathrm{B}}=\rho_{\mathrm{B}}^{eq}$ and $\rho_{\mathrm{AB}}=\rho_{\mathrm{AB}}^{eq}$, and since $K_{r}^{eq}=(\rho_{\mathrm{A}}^{eq}/M_{\mathrm{A}})(\rho_{\mathrm{B}}^{eq}/M_{\mathrm{B}})/(\rho_{\mathrm{AB}}^{eq}/M_{\mathrm{AB}})$, one ends up with $z_{r}=1$. (3) Finally, when the gas is close to “fully dissociated”, $\rho_{\mathrm{AB}}\rightarrow 0$, therefore $z_{r}\rightarrow\infty$.
@@
Downstream of a strong shock, where fully recombined freestream gas experiences sudden heating, this nonequilibrium factor will start out at $z_{r}=0$. At such conditions the non-Boltzmann factors should take on the previously found values $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ to accurately reproduce the QSS dissociation and vibrational energy removal rates. As the gas is advected downstream and relaxes toward post-shock thermal and chemical equilibrium, $z_{r}$ will gradually approach unity. At this point the variable non-Boltzmann factors should both approach unity themselves, so as to reproduce the correct equilibrium reaction and vibrational energy removal rates. Finally, in boundary layers forming near cold surfaces, or in nozzle expansion flows of partially dissociated air, approach to chemical equilibrium can occur from the opposite direction, starting out from some value $z_{r}>1$ and gradually approaching unity as atoms recombine. At present we do not know which precise values the non-Boltzmann factors should take on under such conditions.
Recent three-body QCT [54], Molecular Dynamics-based [55] and state-resolved master equation studies [56] all suggest that recombination preferentially produces diatoms with vibrational energies close to the recombining species’ dissociation threshold, causing over-population in the high-energy tail of the diatom’s vibrational distribution. This in turn potentially leads to non-Boltzmann factors greater than unity in a recombination-dominated gas. However, at this time we do not have first-principles results for recombining flows. Therefore, we choose to cap both non-Boltzmann factors at unity for this recombining regime and propose a linear interpolation function that spans all three limiting cases:
@@
$f_{k/\varepsilon}^{\mathrm{VNB}}(z)=\left\{\begin{array}[]{l l}(1-f_{k/\varepsilon,\,0}^{\mathrm{NB}})\,z+f_{k/\varepsilon,\,0}^{\mathrm{NB}}&\text{if}\quad 0<z<1\\ 1&\text{if}\quad z\geq 1\end{array}\right.$ (28)
@@
This functional form is applied to both non-Boltzmann factors, with $f_{k,0}^{\mathrm{NB}}=0.5$ for the rate coefficient and $f_{\varepsilon,0}^{\mathrm{NB}}=0.85$ for vibrational energy removed by dissociation. Figure 11 illustrates the behavior just described: solid red and blue lines represent the variation of $f_{k}^{\mathrm{VNB}}(z)$ and $f_{\varepsilon}^{\mathrm{VNB}}(z)$ respectively. It should be emphasized that Eq. (28) is ultimately a modeling choice. Linear interpolation was chosen for no other reason than its simplicity. Indeed, an alternative non-linear function has been proposed for $f_{k}^{\mathrm{VNB}}(z)$ (see Eq. (12) of Ref. [57]) and is shown as the black solid line in Fig. 11. Ultimately, the non-Boltzmann factors account for the net effect on rate coefficients and internal energy change in dissociation reactions due to either depleted, or overpopulated high-energy vibrational levels. As discussed in Ref. [53] the degree of overpopulation (when $f_{k/\varepsilon}^{\mathrm{VNB}}(z>1)>1$) can be estimated using depletion statistics combined with state-resolved analysis. However, such estimates have not been verified with first-principles methods, which is why in the MMT model we restrict $f_{k/\varepsilon}^{\mathrm{VNB}}(z)\leq 1$. Therefore, MMT does not currently model overpopulation effects, although it could be extended to include them in the future.
@@
7 Verification against Direct Molecular Simulations
The MMT model presented in this article is a predictive model with a functional form and parametrization based on detailed ab initio results, in contrast to empirical models fit to experimental data. The preceding sections have summarized the numerical values required to evaluate Landau-Teller type vibrational relaxation times and kinetic rate parameters for this simple two-temperature model, incorporating detailed ab inito results for nitrogen and oxygen dissociation [13, 14, 15, 18]. In this section we verify that CFD simulations using MMT accurately reproduce DMS calculations of chemically reacting diatom/atom mixtures which include all physics (i.e. trans-rot.-vib-chemistry relaxation including non-Boltzmann effects). As discussed previously, such CFD comparisons must be performed consistently by including only collision-pair interactions corresponding to the available PESs used in the DMS calculations. This also means that the thermodynamic properties employed in our CFD-MMT verification calculations explicitly do not account for contributions of molecular or atomic species’ electronic excited states. Refer to Tables 2 and 3 for the parameters used to generate the MMT results presented in this section. Further note that constant non-Boltzmann factors of $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ are employed in Eqs. (10) and (17) respectively throughout all MMT calculations presented in this section.

### v2

@@ 6 Approach to chemical equilibrium
$\mathcal{R}_{\mathrm{diss}}-\mathcal{R}_{\mathrm{rec}}]^{\star}=k_{\mathrm{diss}}^{\star}(T)\left(\frac{\rho_{\mathrm{AB}}^{\star}}{M_{\mathrm{AB}}}\right)\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)-k_{\mathrm{rec}}^{\star}(T)\left(\frac{\rho_{\mathrm{A}}^{\star}}{M_{\mathrm{A}}}\right)\left(\frac{\rho_{\mathrm{B}}^{\star}}{M_{\mathrm{B}}}\right)\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)\overset{!}{=}0$ (27)
Thus, from the second equality in Eq. (26) one derives that $k_{\mathrm{rec}}^{\star}(T)=k_{\mathrm{diss}}^{\star}(T)/K_{\mathrm{c}}^{\mathrm{eq}}(T)$, where $k_{\mathrm{diss}}^{\star}(T)$ represents the thermal dissociation rate coefficient. As discussed in Sec. 4.1, we have fit the thermal rate coefficients of all dissociation reactions in Table 2 to the modified Arrhenius form. Thus, in our notation, the equality $k_{\mathrm{rec}}^{\star}(T)=k^{\mathrm{Arr}}(T)/K_{\mathrm{c}}^{\mathrm{eq}}(T)$ must hold.
This indirect way of calculating the recombination rate coefficient is only strictly valid at thermodynamic equilibrium, a state characterized by Maxwell-Boltzmann energy distributions for all species at an unequivocal single temperature $T$. However, to our knowledge all current CFD implementations of two-temperature models invoke the detailed balance relation even when the gas is locally out of thermal equilibrium. One thus assumes approximately that $k_{\mathrm{rec}}=k_{\mathrm{diss}}/K_{\mathrm{c}}^{\mathrm{eq}}(T)$, even when locally nonequilibrium effects may be present.
In the current implementation we also follow this path. We calculate the forward rate by evaluating the dissociation rate coefficient from Eq. (10), yielding $\mathcal{R}_{\mathrm{diss}}^{\mathrm{MMT}}=f_{k}^{\mathrm{NB}}\,Z(T,T_{\mathrm{v}})\,k^{\mathrm{Arr}}\left(\rho_{\mathrm{AB}}/M_{\mathrm{AB}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$. But when we determine the recombination rate coefficient using the approximate detailed balance relation, we explicitly set the nonequilibrium factor $Z$ in Eq. (10) to unity. We do this, because for the recombination rate coefficient a dependence on vibrational temperature is not justified. Recall that the factor $Z(T,T_{\mathrm{v}})$ was introduced in the Marrone-Treanor model to account for vibrational bias found in dissociation and is thus only meaningful for the diatomic reactants possessing vibrational energy modes. Introducing such a bias makes no sense for the recombining atoms. At the same time, the non-Boltzmann factor remains part of the backward rate coefficient expression. We thus obtain $k_{\mathrm{rec}}=f_{k}^{\mathrm{NB}}\,k^{\mathrm{Arr}}(T)/K_{\mathrm{c}}^{\mathrm{eq}}(T)$ and the recombination rate then becomes $\mathcal{R}_{\mathrm{rec}}^{\mathrm{MMT}}=f_{k}^{\mathrm{NB}}\,k^{\mathrm{Arr}}(T)\left(\rho_{A}/M_{A}\right)\left(\rho_{B}/M_{B}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$. When put together, these two expressions yield for the net rate:
$\mathcal{R}_{\mathrm{diss}}-\mathcal{R}_{\mathrm{rec}}=f_{k}^{\mathrm{NB}}\,k^{\mathrm{Arr}}(T)\,\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)\left[Z(T,T_{\mathrm{v}})\left(\frac{\rho_{\mathrm{AB}}}{M_{\mathrm{AB}}}\right)-\frac{1}{K_{\mathrm{c}}^{\mathrm{eq}}(T)}\,\left(\frac{\rho_{\mathrm{A}}}{M_{\mathrm{A}}}\right)\left(\frac{\rho_{\mathrm{B}}}{M_{\mathrm{B}}}\right)\right].$ (28)
This expression makes it clear that, according to the MMT model as formulated in Sec. 2.2, the non-Boltzmann factor re-scales both forward and backward rates simultaneously. This does not constitute a major problem at the location where fully recombined freestream gas initially traverses a shock front. Under such conditions the precise value of the recombination rate will be negligible anyway and only the forward reaction will contribute to any significant degree to Eq. (28). During this early QSS dissociation phase the MMT model will thus reproduce the dissociation rates predicted from an equivalent DMS calculation fairly well.
However, as the gas then proceeds downstream of the shock front, vibrational and translation-rotational modes should gradually relax to a common temperature and eventually the mixture should also approach chemical equilibrium. Forward and backward reactions should then approach their respective thermal rates, i.e. $\mathcal{R}_{\mathrm{diss}}^{\star}=k^{\mathrm{Arr}}(T)$ $\left(\rho_{\mathrm{AB}}^{\star}/M_{\mathrm{AB}}\right)$ $\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ and $\mathcal{R}_{\mathrm{rec}}^{\star}=k^{\mathrm{Arr}}(T)/K_{\mathrm{c}}^{\mathrm{eq}}\left(\rho_{\mathrm{A}}^{\star}/M_{\mathrm{A}}\right)\left(\rho_{\mathrm{B}}^{\star}/M_{\mathrm{B}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ respectively. However, the non-Boltzmann factor originally introduced to help adjust the value of the dissociation rate coefficient for the QSS regime, remains part of the rate coefficient formulation and, if kept at its original value of $f_{k}^{\mathrm{NB}}=0.5$, now prevents the forward and backward processes from reaching their correct thermal rates.
@@
This issue also affects the vibrational energy-chemistry coupling term. Evaluating Eq. (8) and (17) together with Eq. (28) for the present reaction yields:
@@
$w_{\mathrm{v},\mathrm{diss-rec},\mathrm{AB}}^{\mathrm{chem}}=-f_{k}^{\mathrm{NB}}\,f_{\varepsilon}^{\mathrm{NB}}\,k^{\mathrm{Arr}}(T)\,\left(\frac{\rho_{\mathrm{M}}}{M_{\mathrm{M}}}\right)\left[Z(T,T_{\mathrm{v}})\left(\frac{\rho_{\mathrm{AB}}}{M_{\mathrm{AB}}}\right)-\frac{1}{K_{\mathrm{c}}^{\mathrm{eq}}(T)}\,\left(\frac{\rho_{\mathrm{A}}}{M_{\mathrm{A}}}\right)\left(\frac{\rho_{\mathrm{B}}}{M_{\mathrm{B}}}\right)\right]\,\langle\varepsilon_{\mathrm{v},\mathrm{AB}}\rangle_{r}^{\mathrm{Knab}}(T,T_{\mathrm{v}}),$ (29)
@@
with both non-Boltzmann factors seen to be re-scaling the coupling source term. As was the case for the reaction rates themselves, this does not constitute a problem during the early QSS dissociation phase, where only the term for the forward reaction contributes meaningfully to $w_{\mathrm{v},\mathrm{diss-rec},\mathrm{AB}}^{\mathrm{chem}}$. However, when the same constant non-Boltzmann factors are employed regardless of the local conditions, they will continue to incorrectly re-scale both forward and reverse vibrational energy-chemistry removal/replenishment rates even as the gas is meant to approach thermal and chemical equilibrium.
@@
To address this problem, we propose variable non-Boltzmann factors $f_{k}^{\mathrm{VNB}}$ and $f_{\varepsilon}^{\mathrm{VNB}}$, whose values change dynamically with the local state of the gas. Following the approach of Singh and Schwartzentruber [63] we define the reaction’s nonequilibrium concentration ratio:
$\zeta=\frac{(\rho_{\mathrm{A}}/M_{\mathrm{A}})\,(\rho_{\mathrm{B}}/M_{\mathrm{B}})}{(\rho_{\mathrm{AB}}/M_{\mathrm{AB}})\,K_{\mathrm{c}}^{\mathrm{eq}}(T)},$ (30)
@@
which will depend on the local mixture composition and the equilibrium constant for the reaction at temperature $T$. Three limiting cases are of particular interest. (1) When the gas is close to “fully recombined”, i.e. $\rho_{\mathrm{A}}\rightarrow 0$ and $\rho_{\mathrm{B}}\rightarrow 0$, so $\zeta\rightarrow 0$. (2) When the gas takes on its equilibrium composition $\rho_{\mathrm{A}}=\rho_{\mathrm{A}}^{\star}$, $\rho_{\mathrm{B}}=\rho_{\mathrm{B}}^{\star}$ and $\rho_{\mathrm{AB}}=\rho_{\mathrm{AB}}^{\star}$, and since $K_{\mathrm{c}}^{\mathrm{eq}}=(\rho_{\mathrm{A}}^{\star}/M_{\mathrm{A}})(\rho_{\mathrm{B}}^{\star}/M_{\mathrm{B}})/(\rho_{\mathrm{AB}}^{\star}/M_{\mathrm{AB}})$, one ends up with $\zeta=1$. (3) Finally, when the gas is close to “fully dissociated”, $\rho_{\mathrm{AB}}\rightarrow 0$, therefore $\zeta\rightarrow\infty$.
@@
Downstream of a strong shock, where fully recombined freestream gas experiences sudden heating, this concentration ratio will start out at $\zeta=0$. At such conditions the non-Boltzmann factors should take on the previously found values $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ to accurately reproduce the QSS dissociation and vibrational energy removal rates. As the gas is advected downstream and relaxes toward post-shock thermal and chemical equilibrium, $\zeta$ will gradually approach unity. At this point the variable non-Boltzmann factors should both approach unity themselves, so as to reproduce the correct equilibrium reaction and vibrational energy removal rates. Finally, in boundary layers forming near cold surfaces, or in nozzle expansion flows of partially dissociated air, approach to chemical equilibrium will occur from the opposite direction, starting out from some concentration ratio $\zeta>1$ and gradually approaching $\zeta=1$ as atoms recombine. At present we do not know which precise values the two non-Boltzmann factors should take on under such conditions, but there are grounds to expect them both to be greater than unity in a recombination-dominated gas.
The reasoning is as follows. In net-recombining flows where dissociation is negligible, preferential formation of molecules possessing internal energies close to the dissociation threshold $D_{0}$ (a.k.a. preferential recombination) may cause temporary over-population in the high-energy tail of the diatom’s vibrational energy distribution. Such a sudden over-abundance of highly excited diatoms, themselves prime candidates for dissociation, in turn may lead to non-Boltzmann correction factors greater than unity. Note that preferential recombination, the counterpart to preferential dissociation, has been observed in recent three-body QCT [61] and Molecular Dynamics-based [64] studies, as well as a recent state-resolved master equation study [65]. Indeed, in state-specific master equation modeling, preferential recombination into rovibrational levels close to $D_{0}$ is a necessary consequence of enforcing detailed balance relations for state-specific dissociation-recombination reaction pairs (e.g. see Eq. (11) in Ref. [27] or the equivalent Eq. (5) in Ref. [26]).
At the present time though, we do not possess any suitable data to systematically confirm or reject the aforementioned speculations. Studying net-recombining flows from first principles would require us to perform DMS calculations with recombination reactions included, which in turn would presuppose the availability of robust algorithms to simulate on-the-fly three-body trajectories as part of our DMS collision routines. Absent these prerequisites, we are currently unable to estimate the non-Boltzmann factors for concentration rations $\zeta>1$. And as a consequence, we choose to cap both of them at unity for the recombining regime and propose a piece-wise linear interpolation function that spans all three limiting cases:
@@
$f_{k/\varepsilon}^{\mathrm{VNB}}(\zeta)=\left\{\begin{array}[]{l l}(1-f_{k/\varepsilon,\,0}^{\mathrm{NB}})\,\zeta+f_{k/\varepsilon,\,0}^{\mathrm{NB}}&\text{if}\quad 0<\zeta<1\\ 1&\text{if}\quad\zeta\geq 1\end{array}\right$. (31)
@@
This functional form is applied to both non-Boltzmann factors, with $f_{k,0}^{\mathrm{NB}}=0.5$ for the rate coefficient and $f_{\varepsilon,0}^{\mathrm{NB}}=0.85$ for vibrational energy removed by dissociation. Figure 13 illustrates the behavior just described. Solid red and blue lines represent the variation of $f_{k}^{\mathrm{VNB}}(\zeta)$ and $f_{\varepsilon}^{\mathrm{VNB}}(\zeta)$ respectively. It should be emphasized that Eq. (31) is ultimately a modeling choice. Linear interpolation was chosen for no other reason than its simplicity. Indeed, an alternative non-linear function has been proposed for $f_{k}^{\mathrm{VNB}}(\zeta)$ (see Eq. (12) of Ref. [66]) and is shown as the black solid line in Fig. 13. Ultimately, the non-Boltzmann factors account for the net effect on rate coefficients and internal energy change in dissociation reactions due to either depleted, or overpopulated high-energy vibrational levels. As discussed in Ref. [63] the degree of overpopulation (when $f_{k/\varepsilon}^{\mathrm{VNB}}(\zeta>1)>1$) can be estimated using depletion statistics combined with state-resolved analysis. However, such estimates have so far not been thoroughly confirmed by first-principles methods, which is why in the MMT model we restrict $f_{k/\varepsilon}^{\mathrm{VNB}}(\zeta)\leq 1$. Therefore, MMT does not currently model overpopulation effects, although it could be extended to include them in the future.
@@
7 Benchmarking against Direct Molecular Simulations
Our overarching goal for the Modified Marrone-Treanor model has been to construct a predictive dissociation model with a functional form and parametrization based as much as possible on detailed ab initio results, in contrast to existing empirical models that have primarily been fit to experimental data. In the preceding sections we have presented the MMT model’s analytical formulation and listed the parameters required to evaluate dissociation rate coefficients, vibrational energy removal rates, as well as characteristic vibrational relaxation times, all derived from detailed ab initio calculations for nitrogen and oxygen [15, 21, 22, 25, 37].
A complete validation of MMT for real-world applications will require large-scale CFD simulations of nonequilibrium reacting flows produced at hypersonic flight conditions, or in ground test facilities. However, before achieving this milestone, we must assess to what degree our CFD implementation of the MMT model agrees with DMS results that already describe at the microscopic scale the most relevant thermochemical nonequilibrium processes in $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$) mixtures. In these comparison calculations we focus on the dissociation chemistry coupled to vibrational relaxation, which in the CFD model is fully driven by the source terms of Eqs. (1) and (5). It is most convenient to study these phenomena isolated from other fluid dynamics effects, such as convection and diffusion, by simulating simple space-homogeneous heat baths (also referred to as 0D reactors), either under isothermal (constant translational temperature), or adiabatic (constant total energy) conditions.
In order to achieve the most meaningful comparison, we must take care to only include physical processes in the CFD-MMT calculations that are also captured by DMS. Thus, the chemistry parameters taken from Table 2 (either sub-tables) and vibrational relaxation from Table 3 were derived on the same PES set as the one employed in the DMS reference calculations. This implies that in both solution methods the dissociation and vibrational relaxation processes are based exclusively on electronically adiabatic trajectories between ground-state atoms and molecules. It further means that the thermodynamic properties (i.e. enthalpies, heat capacities, equilibrium constants, etc.) employed in our CFD-MMT benchmarking calculations explicitly do not account for contributions of molecular or atomic species’ electronic excited states.
Refer to Tables 2 and 3 for the parameters used to generate the MMT results presented in this section. Further note that constant non-Boltzmann factors of $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ are employed in Eqs. (10) and (17) respectively throughout all MMT calculations presented in this section.